OpenStudy (anonymous):

PLEASE HELP ASAP. How much greater is the volume of a cube if the length of each side is multiplied by 10?

OpenStudy (anonymous):

Say your cube has side length 4, then 4*4*4 = 64, but if you multiply each side by 10 then it becomes: 40*40*40=64000, which is a 99900% increase.

OpenStudy (anonymous):

Thank you for helping me, If the side lengths are not given, how much greater would the volume be?

OpenStudy (zehanz):

Original length of side is a. Volume is a*a*a=a³ New side is 10a, Volume = 10a*10a*10a=... Can you see where this is getting to? BTW, from the question I wouldn't have guessed giving a percentage. I'd say a multiplication factor will do.

OpenStudy (anonymous):

It would still always be 99900% increase, or just multiply the original volume by 1000 to get new volume where sides are multiplied by 10.

OpenStudy (zehanz):

@Zelda is right! 10a*10a*10a = 10*10*10*a*a*a = 1000a³, so that is 1000 times the original value.